OpenStudy (anonymous):

Identify the solution to the equation 2 |3x -5| = 8

OpenStudy (jiteshmeghwal9):

first of all u must remember that if there is a unknown value in the mod then the value will come both in positive or negative.

OpenStudy (anonymous):

So I would leave the 2 out of the equation until I solve for X in the absolute value?

OpenStudy (jiteshmeghwal9):

so, case I:2(3x-5)=8 6x-10=8 6x=8+10 6x=18 x=18/6=3

OpenStudy (jiteshmeghwal9):

case II:2(-(3x-5))=8 2(-3x+5)=8 -6x+10=8 -6x=8-10 -6x=-2 x=-2/-6=1/3

OpenStudy (jiteshmeghwal9):

hence,x=3 & 1/3.

OpenStudy (jiteshmeghwal9):

gt it @iAmPureSwagger ???

OpenStudy (anonymous):

Oh my gosh one little mistake on my paper and I messed up the entire equation! @jiteshmeghwal9 Thanks for your help, and yes you helped clarify!

OpenStudy (jiteshmeghwal9):

yw:) Nice to help u @iAmPureSwagger :)